# Composite Flow Matching for Reinforcement Learning with Shifted-Dynamics Data

### v2

@@ Appendix D Algorithm of CompFlow built on Soft-Actor-Critic
For the policy $\pi$, we then decompose the performance gap
$\underbrace{\eta_{\mathcal{M}_{\text{tar}}}\!\bigl{(}\pi_{\text{tar}}^{*}\bigr{)}-\eta_{\mathcal{M}_{\text{tar}}}\!\bigl{(}\pi\bigr{)}}_{\text{desired gap}}$
$=\;\underbrace{\bigl{[}\eta_{\mathcal{M}_{\text{tar}}}(\pi_{\text{tar}}^{*})-\eta_{\mathcal{M}_{\text{src}}}(\pi_{\text{tar}}^{*})\bigr{]}}_{\text{model gap (a)}}+\;\underbrace{\bigl{[}\eta_{\mathcal{M}_{\text{src}}}(\pi_{\text{tar}}^{*})-\eta_{\mathcal{M}_{\text{src}}}(\pi)\bigr{]}}_{\text{learning gap (b)}}+\;\underbrace{\bigl{[}\eta_{\mathcal{M}_{\text{src}}}(\pi)-\eta_{\mathcal{M}_{\text{tar}}}(\pi)\bigr{]}}_{\text{model gap (c)}}.$ (15)
Let’s call the behavior cloning policy on the offline source dataset $\pi_{\text{bc}}$, and we use $\varepsilon_{\text{bc}}:=\eta_{\mathcal{M}_{\text{src}}}(\pi_{\text{tar}}^{*})-\eta_{\mathcal{M}_{\text{src}}}(\pi_{\text{bc}})$ to denote the learning gap, i.e., the (b) term in decomposition  G.
$\eta_{\mathcal{M}_{\text{tar}}}\!\bigl{(}\pi_{\text{tar}}^{*}\bigr{)}-\eta_{\mathcal{M}_{\text{tar}}}\!\bigl{(}\hat{\pi}_{\text{bc}}\bigr{)}$ $\leq\frac{4L_{r}(1+\gamma)}{(1-\gamma)(1-\gamma L_{p})}\Delta_{W_{2}}\;+\;\varepsilon_{\text{bc}}.$ (16)
During the online learning stage, we collect samples with large dynamics gap. Suppose in our policy $\hat{\pi}$, we have all samples with $\Delta_{W_{2}}>\kappa$ being replaced with samples from the target MDP. Considering MSE errors introduced by flow matching, as in the proof of Theorem 3.1, the errors could be
$Err_{\kappa}\leq\text{Var}(s^{\prime}|\nu)_{\text{tar}}+O\left(\frac{\sigma^{2}_{\text{tar}}}{\sqrt{N_{\text{tar}}}}\right)$ (17)
We denote the source and target distribution of flow matching by $p_{\text{src}}$ and $p_{\text{tar}}$, and the pushforward distribution generated by flow matching as $\hat{p}_{\text{tar}}$. By Lemma  G.1, we can bound the Wasserstein-2 distance estimation error by
@@
$\left|W_{2}(p_{\mathrm{src}},p_{\mathrm{tar}})-W_{2}(p_{\mathrm{src}},\hat{p}_{\mathrm{tar}})\right|\leq\sqrt{Err_{\kappa}}$
Let’s use $\varepsilon_{N}:=\eta_{\mathcal{M}_{\text{src}}}(\pi_{\text{tar}}^{*})-\eta_{\mathcal{M}_{\text{src}}}(\hat{\pi})$ to denote the learning gap of the fine-tuned policy $\hat{\pi}$ in decomposition  G. Therefore, the performance gap of $\hat{\pi}$ is:
$\eta_{\mathcal{M}_{\text{tar}}}\!\bigl{(}\pi_{\text{tar}}^{*}\bigr{)}-\eta_{\mathcal{M}_{\text{tar}}}\!\bigl{(}\hat{\pi}\bigr{)}$
$\leq\frac{4L_{r}(1+\gamma)}{(1-\gamma)(1-\gamma L_{p})}\,\left(\kappa+\sqrt{\text{Var}(s^{\prime}|\nu)_{\text{tar}}+O\left(\frac{\sigma^{2}_{\text{tar}}}{\sqrt{N_{\text{tar}}}}\right)}\right)\;+\;\varepsilon_{N}.$ (18)
We assume that the model has been trained for sufficiently long time and the network capacity is sufficient so $\varepsilon_{N}$ and $\varepsilon_{\text{bc}}$ are small and have negligible influences. Therefore, by subtracting RHS of Equation G from RHS of Equation 16, we obtain the performance gap reduces at least by
$\frac{4L_{r}(1+\gamma)}{(1-\gamma)(1-\gamma L_{p})}\left(\Delta_{W_{2}}-\kappa-\sqrt{\text{Var}(s^{\prime}|\nu)_{\text{tar}}+O\left(\frac{\sigma^{2}_{\text{tar}}}{\sqrt{N_{\text{tar}}}}\right)}\right).$ (19)
Appendix H Proof of Proposition 3.2
@@
Proposition H.1 (Shared-Latent Coupling Approximates $W_{2}$-Optimal Transport).
@@
Notation. Fix a state–action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$, and omit $(s,a)$ in the following for notational simplicity. Let the latent distribution be $p_{0}(x_{0})=\mathcal{N}(0,\mathbf{I})$. Define the source transition distribution as $p_{1}(x_{1}):=p_{\rm src}(s^{\prime}|s,a)$, and the target transition distribution as $p_{2}(x_{2}):=p_{\text{tar}}(s^{\prime}|s,a)$.
@@
According to Proposition 3.1 in [17], as $\eta\to\infty$, mass transport in the conditional variables $(s,a)$ vanishes. That is, the transport is concentrated exclusively in the next-state space, and optimal transport is performed independently within each fixed $(s,a)$ pair when $\eta\rightarrow\infty$.
Therefore, for any fixed $(s,a)$, we can invoke Theorem 4.2 from [42], which guarantees that the expected transport cost under the batch-trained flow $\psi^{k}_{\phi}$ converges to the squared 2-Wasserstein distance between $p_{1}$ and $p_{2}$ as the batch size $k\to\infty$. Hence,
@@
Appendix I Additional Experimental Results
I.1 Additional Main Results
We provide additional results on the Walker2D Kinematic and Frictional shift tasks in Table 2. Across these 12 additional tasks, CompFlow consistently outperforms other methods—achieving the best performance on 10 tasks, tying on 2, and losing on only 1.
Dataset Task Name BC-SAC H2O BC-VGDF BC-PAR Ours
MR Walker2D (Kinematic) $2973\pm 185$ $1967\pm 851$ $1586\pm 923$ $948\pm 131$ $1568\pm 1315$
MR Halfcheetah (Friction) $7799\pm 157$ $6397\pm 673$ $7829\pm 821$ $8056\pm 512$ $8241\pm 180$
MR Hopper (Friction) $228\pm 2$ $229\pm 1$ $230\pm 3$ $232\pm 5$ $280\pm 27$
MR Walker2D (Friction) $311\pm 5$ $296\pm 14$ $302\pm 12$ $321\pm 26$ $344\pm 20$
M Walker2D (Kinematic) $1966\pm 1155$ $1965\pm 568$ $1921\pm 928$ $806\pm 278$ $2039\pm 936$
M Halfcheetah (Friction) $7108\pm 1001$ $6968\pm 846$ $6802\pm 956$ $7800\pm 525$ $7871\pm 238$
M Hopper (Friction) $232\pm 1$ $228\pm 4$ $229\pm 4$ $233\pm 5$ $300\pm 66$
M Walker2D (Friction) $286\pm 54$ $298\pm 45$ $289\pm 47$ $308\pm 24$ $320\pm 31$
ME Walker2D (Kinematic) $850\pm 953$ $1514\pm 782$ $1204\pm 734$ $755\pm 268$ $1511\pm 1206$
ME Halfcheetah (Friction) $4185\pm 732$ $2140\pm 733$ $4078\pm 1032$ $4989\pm 500$ $5596\pm 1557$
ME Hopper (Friction) $230\pm 3$ $232\pm 5$ $230\pm 6$ $232\pm 8$ $266\pm 70$
ME Walker2D (Friction) $240\pm 114$ $258\pm 8$ $245\pm 51$ $242\pm 24$ $326\pm 26$
Average Return $2201$ $1874$ $2079$ $2077$ $2389$
Table 2: Comparison of policy performance across frictional and kinematic dynamics shifts under three dataset regimes: MR = Medium Replay, M = Medium, ME = Medium Expert. Green highlights indicate the top-performing method with $\geq$2improvement; yellow indicates ties within 2%.
Appendix J Experimental Details
@@ I.1 Additional Main Results
J.1 Environment Setting
J.1.1 Source Domain Dataset
We use the MuJoCo datasets from D4RL [16] as our source domain data. These datasets are collected from continuous control environments in Gym [4], simulated using the MuJoCo physics engine [56]. We focus on three benchmark tasks: HalfCheetah, Hopper, and Walker2d, and evaluate across three dataset types: medium, medium-replay, and medium-expert.
@@
J.1.2 Kinematic Shift Tasks
We use Kinematic Shift Tasks from the benchmark [33]. We select most shift level ’hard’ to make the tasks more challenging
@@
$\bullet$ Hopper Kinematic Shift: the rotation range of the foot joint is modified from $[-45,45]$ to $[-9,9]$:
@@
$\bullet$ Walker2D Kinematic Shift: the rotation range of the foot joint is modified from $[-45,45]$ to $[-9,9]$:
@@
J.1.3 Morphology Shift Tasks
We use Morphology Shift Tasks from the benchmark [33]. We select most shift level ’hard’ to make the tasks more challenging
@@
J.1.4 Friction Shift Tasks
Following [33], the friction shift is implemented by altering the $\mathsf{friction}$ attribute in the $\mathsf{geom}$ elements. The frictional components are adjusted to 5.0 times the frictional components in the source domain. The following is an example for the Hopper robot.
@@
J.2 Implementation Details
BC-SAC: This baseline leverages both offline source domain data and online target domain transitions for policy learning. Since learning from offline data requires conservatism while online data does not, we incorporate a behavior cloning term into the actor update of the SAC algorithm. Specifically, the critic is updated using standard Bellman loss on the combined offline and online datasets, and the actor is optimized as:
$\mathcal{L}_{\text{actor}}=\lambda\cdot\mathbb{E}_{s\sim\mathcal{D}_{\text{src}}\cup\mathcal{D}_{\text{tar}},\ a\sim\pi_{\varphi}(\cdot|s)}\left[\min_{i=1,2}Q_{\varsigma_{i}}(s,a)-\alpha\log\pi_{\varphi}(\cdot|s)\right]+\mathbb{E}_{(s,a)\sim\mathcal{D}_{\text{src}},\ \hat{a}\sim\pi_{\varphi}(\cdot|s)}\left[(a-\hat{a})^{2}\right],$ (20)
where $\lambda=\frac{\omega}{\frac{1}{N}\sum_{(s_{j},a_{j})}\min_{i=1,2}Q_{\varsigma_{i}}(s_{j},a_{j})}$ and $\omega\in\mathbb{R}^{+}$ is a normalization coefficient. We train BC-SAC for 400K gradient steps, collecting target domain data every 10 steps. We use the hyperparameters recommended in [33].
H2O: H2O [41] trains domain classifiers to estimate dynamics gaps and uses them as importance sampling weights during critic training. It also incorporates a CQL loss to encourage conservatism. Since the original H2O is designed for the Online-Offline setting (online source, offline target), we adapt the objective to the Offline-Online setting. The critic loss is:
$\mathcal{L}_{\text{critic}}=$ $\ \mathbb{E}_{\mathcal{D}_{\text{tar}}}\left[(Q_{\varsigma_{i}}(s,a)-y)^{2}\right]+\mathbb{E}_{\mathcal{D}_{\text{src}}}\left[\omega(s,a,s^{\prime})(Q_{\varsigma_{i}}(s,a)-y)^{2}\right]$
$+\beta_{\text{CQL}}\left(\mathbb{E}_{s\sim\mathcal{D}_{\text{src}},\ \hat{a}\sim\pi_{\varphi}(\cdot|s)}\left[\omega(s,a,s^{\prime})Q_{\varsigma_{i}}(s,\hat{a})\right]-\mathbb{E}_{\mathcal{D}_{\text{src}}}\left[\omega(s,a,s^{\prime})Q_{\varsigma_{i}}(s,a)\right]\right),\quad i\in\{1,2\},$ (21)
@@
BC-VGDF: BC-VGDF [59] filters source transitions whose estimated values align closely with those from the target domain. It trains an ensemble of dynamics models to predict next states from raw state-action pairs under the target dynamics. Each predicted next state is evaluated by the policy to obtain a value ensemble $\{Q(s^{\prime}_{i},a^{\prime}_{i})\}_{i=1}^{M}$, forming a Gaussian distribution. A fixed percentage ($\xi\%$) of source transitions with the highest likelihood under this distribution are retained. The critic loss is:
$\mathcal{L}_{\text{critic}}=$ $\ \mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}_{\text{tar}}}\left[\left(Q_{\varsigma_{i}}(s,a)-y\right)^{2}\right]$
$+\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}_{\text{src}}}\left[\mathbf{1}\left(\Lambda(s,a,s^{\prime})>\Lambda_{\xi\%}\right)\left(Q_{\varsigma_{i}}(s,a)-y\right)^{2}\right],\quad i\in\{1,2\},$ (22)

### v3

@@ Appendix E Algorithm of CompFlow built on Soft-Actor-Critic
Step B: FM-based uniform $W_{2}$ estimation bound (construction of $\alpha$).
We need a high-probability uniform control of the error of the rectified-flow $W_{2}$ estimator to ensure that thresholding by $\kappa$ actually caps the on-policy mismatch for $\hat{\pi}$.
B.1. Flow-to-map error. Let $T$ be the Brenier OT map from $\hat{p}_{\rm off}(\cdot\mid s,a)$ to $p_{\rm on}(\cdot\mid s,a)$ (guaranteed by standard assumptions), and let $\widehat{T}$ be the terminal map obtained by integrating the learned FM velocity field $\hat{v}$ from $t=0$ to $t=1$. Under the linear path and Lipschitz regularity of $v^{\star}$ and $\hat{v}$ (Theorem 3.1 assumptions), a stability argument (Gronwall) yields
$\mathbb{E}_{X_{0}\sim\hat{p}_{\rm off}}\bigl\|\widehat{T}(X_{0})-T(X_{0})\bigr\|_{2}^{2}\;\leq\;K_{\mathrm{stab}}\;\int_{0}^{1}\mathbb{E}\bigl\|\hat{v}(X_{t},t)-v^{\star}(X_{t},t)\bigr\|_{2}^{2}\,dt,$ (27)
for some finite $K_{\mathrm{stab}}$ depending on Lipschitz constants of the flow. By the “master bound” (Eq. (7) in Theorem 3.1), uniformly over $t\in[0,1]$, with prob. $\geq 1-\delta$,
$\mathbb{E}\bigl\|\hat{v}-v^{\star}\bigr\|_{2}^{2}\;\leq\;C_{0}\,\Gamma_{N_{\rm on},\delta}+C_{1}\,W_{2}(\hat{p}_{\rm off},p_{\rm on})\,\Gamma_{N_{\rm on},\delta}.$ (28)
So we have
$\underbrace{\mathbb{E}\|\widehat{T}(X_{0})-T(X_{0})\|_{2}^{2}}_{\text{terminal map MSE}}$ $\overset{\text{Gronwall}}{\leq}K_{\mathrm{stab}}\int_{0}^{1}\underbrace{\mathbb{E}\|\hat{v}-v^{\star}\|_{2}^{2}}_{\text{FM excess risk at }t}dt$ (29)
$\leq K_{\mathrm{stab}}\left(C_{0}+C_{1}\,W_{2}(\hat{p}_{\rm off},p_{\rm on})\right)\Gamma_{N_{\rm on},\delta}$ (30)
B.2. From map MSE to W2 error. By the coupling construction $(\hat{s},s)=(\widehat{T}(X_{0}),T(X_{0}))$ with $X_{0}\sim\hat{p}_{\rm off}$ and the triangle inequality for $W_{2}$,
$\sup_{(s,a)}\;\Bigl|\,\widehat{\Delta}_{W_{2}}(s,a)-\Delta_{W_{2}}(s,a)\,\Bigr|\;\leq\;\sup_{(s,a)}\;W_{2}\bigl(\hat{p}_{\rm on},p_{\rm on}\bigr)$ (31)
Define the coupling $\gamma_{(s,a)}:=(T,\widehat{T})_{\#}\hat{p}_{\rm off}(\cdot\mid s,a)$, i.e., if $S:=T(X_{0})$ and $\widehat{S}:=\widehat{T}(X_{0})$ then $(S,\widehat{S})\sim\gamma_{(s,a)}$ and $\gamma_{(s,a)}\in\Pi\!\big(p_{\rm on}(\cdot\mid s,a),\hat{p}_{\rm on}(\cdot\mid s,a)\big)$. By the definition of the $2$-Wasserstein distance,
$W_{2}^{2}\!\big(p_{\rm on}(\cdot\mid s,a),\hat{p}_{\rm on}(\cdot\mid s,a)\big)$ $=\inf_{\gamma\in\Pi(p_{\rm on},\hat{p}_{\rm on})}\int\!\|x-y\|_{2}^{2}\,\mathrm{d}\gamma(x,y)$ (32)
$\leq\;\int\!\|x-y\|_{2}^{2}\,\mathrm{d}\gamma_{(s,a)}(x,y)=\mathbb{E}\,\|\widehat{T}(X_{0})-T(X_{0})\|_{2}^{2}.$ (33)
Taking square roots gives, for each $(s,a)$,
$W_{2}\!\big(p_{\rm on}(\cdot\mid s,a),\hat{p}_{\rm on}(\cdot\mid s,a)\big)\;\leq\;\sqrt{\mathbb{E}\,\|\widehat{T}(X_{0})-T(X_{0})\|_{2}^{2}}.$
Finally, taking the supremum over $(s,a)$ preserves the inequality:
$\sup_{(s,a)}W_{2}\!\big(\hat{p}_{\rm on}(\cdot\mid s,a),p_{\rm on}(\cdot\mid s,a)\big)\;\leq\;\sup_{(s,a)}\sqrt{\mathbb{E}\,\|\widehat{T}(X_{0})-T(X_{0})\|_{2}^{2}}.$ (34)
Combining (34) and (31) gives
$\sup_{(s,a)}\Bigl|\,\widehat{\Delta}_{W_{2}}(s,a)-\Delta_{W_{2}}(s,a)\,\Bigr|\;\leq\;\alpha_{N_{\rm on},\delta}:=\sqrt{K_{\mathrm{stab}}\left(C_{0}+C_{1}\,W_{2}(\hat{p}_{\rm off},p_{\rm on})\right)\Gamma_{N_{\rm on},\delta}},$
with probability at least $1-\delta$.
B.3. Capping the on-policy gap for $\hat{\pi}$. By construction of the replacement rule, for any $(s,a)$ encountered during fine-tuning of $\hat{\pi}$ we have either $\widehat{\Delta}_{W_{2}}(s,a)\leq\kappa$ (kept offline sample) or we replaced it by an online sample. In either case, on those $(s,a)$ we ensure
$\Delta_{W_{2}}(s,a)\;\leq\;\widehat{\Delta}_{W_{2}}(s,a)+\alpha_{N_{\rm on},\delta}\;\leq\;\kappa+\alpha_{N_{\rm on},\delta}.$
Therefore the on-policy gap of $\hat{\pi}$ obeys
$\Delta_{W_{2}}^{(\hat{\pi})}\;\leq\;\kappa+\alpha_{N_{\rm on},\delta}.$ (35)
Step C: Learning-gap bounds for $\pi_{\mathrm{bc}}$ and $\widehat{\pi}$ (construction of $\beta$).
Recall the three-term decomposition for any policy $\pi$:
$\eta_{\rm on}(\pi_{\rm on}^{\star})-\eta_{\rm on}(\pi)=\underbrace{\bigl(\eta_{\rm on}(\pi_{\rm on}^{\star})-\eta_{\rm off}(\pi_{\rm on}^{\star})\bigr)}_{\text{model(a)}}+\underbrace{\bigl(\eta_{\rm off}(\pi_{\rm on}^{\star})-\eta_{\rm off}(\pi)\bigr)}_{\text{learning(b)}}+\underbrace{\bigl(\eta_{\rm off}(\pi)-\eta_{\rm on}(\pi)\bigr)}_{\text{model(c)}}.$ (36)
We now bound the learning term (b) by ERM generalization.
Policies are learned by ERM on a surrogate imitation loss $\mathcal{L}(\pi)$ over a policy class $\Pi$ with Rademacher complexity $\mathfrak{R}_{N}(\Pi)$. There exists a calibration constant $C_{\mathrm{val}}$ (depends on concentrability/Lipschitzness; fixed for the class) such that, for any $\pi$,
$\eta_{\rm off}(\pi_{\rm on}^{\star})-\eta_{\rm off}(\pi)\;\leq\;C_{\mathrm{val}}\Bigl(\mathcal{L}(\pi)-\inf_{\pi^{\prime}\in\Pi}\mathcal{L}(\pi^{\prime})\Bigr).$ (37)
Moreover, for datasets of sizes $N_{\rm off},N_{\mathrm{mod}}$, with probability at least $1-\delta$,
$\mathcal{L}(\hat{\pi})-\inf_{\pi^{\prime}\in\Pi}\mathcal{L}(\pi^{\prime})\;\leq\;C_{\Pi}\Bigl(\mathfrak{R}_{N}(\Pi)+\sqrt{\tfrac{\log(1/\delta)}{N}}\Bigr),\qquad N\in\{N_{\rm off},N_{\mathrm{mod}}\},$ (38)
for an absolute constant $C_{\Pi}$.
Let $\pi_{\mathrm{bc}}$ be the behavior cloning policy trained on the offline dataset of size $N_{\rm off}$, and let $\hat{\pi}$ be ERM on the modified dataset (size $N_{\mathrm{mod}}$): for any $(s,a)$ encountered, if the FM-estimated gap satisfies $\widehat{\Delta}_{W_{2}}(s,a)>\kappa$, replace the offline $(s,a,s^{\prime})$ sample by an online $(s,a,s^{\prime}_{\rm on})$ sample (collecting $N_{\rm on}$ such online transitions), leaving all others unchanged. Denote $N_{\mathrm{mod}}$ the resulting (modified) dataset size.
By calibration (37) and the generalization inequality (38), with probability at least $1-\delta$ we have
$\eta_{\rm off}(\pi_{\rm on}^{\star})-\eta_{\rm off}(\pi_{\mathrm{bc}})\;\leq\;C_{\mathrm{val}}\;C_{\Pi}\Bigl(\mathfrak{R}_{N_{\rm off}}(\Pi)+\sqrt{\tfrac{\log(1/\delta)}{N_{\rm off}}}\Bigr),$
and similarly
$\eta_{\rm off}(\pi_{\rm on}^{\star})-\eta_{\rm off}(\hat{\pi})\;\leq\;C_{\mathrm{val}}\;C_{\Pi}\Bigl(\mathfrak{R}_{N_{\mathrm{mod}}}(\Pi)+\sqrt{\tfrac{\log(1/\delta)}{N_{\mathrm{mod}}}}\Bigr).$
By a union bound (probability $1-2\delta$) and adding these two upper bounds we obtain a common envelope
$\max\Bigl\{\eta_{\rm off}(\pi_{\rm on}^{\star})-\eta_{\rm off}(\pi_{\mathrm{bc}}),\;\eta_{\rm off}(\pi_{\rm on}^{\star})-\eta_{\rm off}(\hat{\pi})\Bigr\}\;\leq\;\beta,$ (39)
with
@@
$\beta:=C_{\mathrm{val}}C_{\Pi}\!\left(\mathfrak{R}_{N_{\rm off}}(\Pi)+\mathfrak{R}_{N_{\mathrm{mod}}}(\Pi)+\sqrt{\tfrac{\log(2/\delta)}{N_{\rm off}}}+\sqrt{\tfrac{\log(2/\delta)}{N_{\mathrm{mod}}}}\right).$ (40)
Step D: Assemble bounds and compare.
Incorporate (26) and (39) to (36) with $\pi=\pi_{\mathrm{bc}}$, we have with high probability,
$\eta_{\rm on}(\pi_{\rm on}^{\star})-\eta_{\rm on}(\pi_{\mathrm{bc}})\leq\bigl(\eta_{\rm on}(\pi_{\rm on}^{\star})-\eta_{\rm off}(\pi_{\rm on}^{\star})\bigr)+\beta+\frac{2L_{r}(1+\gamma)}{(1-\gamma)(1-\gamma L_{p})}\Delta_{W_{2}}$
Incorporate (26) and (39) to (36) with $\pi=\hat{\pi}$, we have with high probability
$\eta_{\rm on}(\pi_{\rm on}^{\star})-\eta_{\rm on}(\widehat{\pi})\;\leq\;\bigl(\eta_{\rm on}(\pi_{\rm on}^{\star})-\eta_{\rm off}(\pi_{\rm on}^{\star})\bigr)+\beta+\frac{2L_{r}(1+\gamma)}{(1-\gamma)(1-\gamma L_{p})}(\kappa+\alpha_{N_{\rm on},\delta})$
The difference between these two upper bounds is:
$\frac{2L_{r}(1+\gamma)}{(1-\gamma)(1-\gamma L_{p})}\bigl(\Delta_{W_{2}}-\kappa-\alpha_{N_{\rm on},\delta}\bigr).$
Absorbing $K_{\mathrm{stab}}$ into $C_{0}$ and $C_{1}$ gets the result. ∎
Appendix I Proof of Proposition 3.3
@@
Proposition I.1 (Shared-Latent Coupling Approximates $W_{2}$-Optimal Transport).
@@
Notation. Fix a state–action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$, and omit $(s,a)$ in the following for notational simplicity. Let the latent distribution be $p_{0}(x_{0})=\mathcal{N}(0,\mathbf{I})$. Define the offline transition distribution as $p_{1}(x_{1}):=p_{\rm off}(s^{\prime}|s,a)$, and the online transition distribution as $p_{2}(x_{2}):=p_{\rm{on}}(s^{\prime}|s,a)$.
@@
According to Proposition 3.1 in [generale2024conditional], as $\eta\to\infty$, mass transport in the conditional variables $(s,a)$ vanishes. That is, the transport is concentrated exclusively in the next-state space, and optimal transport is performed independently within each fixed $(s,a)$ pair when $\eta\rightarrow\infty$.
Therefore, for any fixed $(s,a)$, we can invoke Theorem 4.2 from [pooladian2023multisample], which guarantees that the expected transport cost under the batch-trained flow $\psi^{k}_{\phi}$ converges to the squared 2-Wasserstein distance between $p_{1}$ and $p_{2}$ as the batch size $k\to\infty$. Hence,
@@
Appendix J Additional Experimental Results
Dataset Task Name SAC BC-SAC H2O BC-VGDF BC-PAR Ours
MR HalfCheetah (Morphology) $1239\pm 107$ $1896\pm 357$ $1294\pm 153$ $1342\pm 252$ $1967\pm 123$ $2267\pm 280$
MR HalfCheetah (Kinematic) $2511\pm 476$ $4275\pm 129$ $3787\pm 205$ $3821\pm 187$ $3751\pm 113$ $4448\pm 227$
M HalfCheetah (Morphology) $078\pm 165 11$$199\pm 184$ $1277\pm 154$ $1014\pm 141$ $1263\pm 163$ $1654\pm 136$
M HalfCheetah (Kinematic) $2331\pm 392$ $4650\pm 169$ $4572\pm 84$ $4207\pm 108$ $4304\pm 231$ $4651\pm 89$
ME HalfCheetah (Morphology) $1115\pm 207$ $1067\pm 208$ $1034\pm 206$ $048\pm 202 11$$196\pm 203$ $1834\pm 193$
ME HalfCheetah (Kinematic) $1843\pm 1185$ $2650\pm 447$ $3739\pm 196$ $3409\pm 207$ $3108\pm 500$ $4537\pm 74$
Table 2: Comparison of return under the extreme difficult settings. MR = Medium Replay, M = Medium, ME = Medium Expert. A cell is green if the method has the highest mean and improves over the second best by at least 2%. Cells within 2% of the top mean are marked in yellow.
To further increase the task difficulty, we manually modified the benchmark configurations to amplify the dynamic differences beyond the given settings. For example, in the morphology task, we drastically reduced thigh segment size—making the legs much shorter, in the kinematic task, we almost immobilized the back-thigh joint by severely limiting its rotation range. As shown in Table 2, our method clearly outperforms all other baselines in 5 out of 6 settings, while achieving ties in the remaining one.
Appendix K Experimental Details of Gym-MuJoCo
@@
K.1 Environment Setting
K.1.1 Offline Dataset
We use the MuJoCo datasets from D4RL [fu2020d4rl] as our offline data. These datasets are collected from continuous control environments in Gym [brockman2016openai], simulated using the MuJoCo physics engine [todorov2012mujoco]. We focus on three benchmark tasks: HalfCheetah, Hopper, and Walker2d, and evaluate across three dataset types: medium, medium-replay, and medium-expert.
@@
K.1.2 Kinematic Shift Tasks
We use Kinematic Shift Tasks from the benchmark [lyu2024odrl]. We select most shift level ’hard’ to make the tasks more challenging
@@
$\bullet$ Hopper Kinematic Shift: the rotation range of the foot joint is modified from [45,45][-45,45] to [9,9][-9,9]:
@@
$\bullet$ Walker2D Kinematic Shift: the rotation range of the foot joint is modified from [45,45][-45,45] to [9,9][-9,9]:
@@
K.1.3 Morphology Shift Tasks
We use Morphology Shift Tasks from the benchmark [lyu2024odrl]. We select most shift level ’hard’ to make the tasks more challenging
@@
K.1.4 Friction Shift Tasks
Following [lyu2024odrl], the friction shift is implemented by altering the $\mathsf{friction}$ attribute in the $\mathsf{geom}$ elements. The frictional components are adjusted to 5.0 times the frictional components in the offline environment. The following is an example for the Hopper robot.
@@
K.2 Implementation Details
BC-SAC: This baseline leverages both offline and online transitions for policy learning. Since learning from offline data requires conservatism while online data does not, we incorporate a behavior cloning term into the actor update of the SAC algorithm. Specifically, the critic is updated using standard Bellman loss on the combined offline and online datasets, and the actor is optimized as:
$\mathcal{L}_{\text{actor}}=\lambda\cdot\mathbb{E}_{s\sim\mathcal{D}_{\text{off}}\cup\mathcal{D}_{\text{on}},\ a\sim\pi_{\varphi}(\cdot|s)}\left[\min_{i=1,2}Q_{\varsigma_{i}}(s,a)-\alpha\log\pi_{\varphi}(\cdot|s)\right]+\mathbb{E}_{(s,a)\sim\mathcal{D}_{\text{off}},\ \hat{a}\sim\pi_{\varphi}(\cdot|s)}\left[(a-\hat{a})^{2}\right],$ (41)
where $\lambda=\frac{\omega}{\frac{1}{N}\sum_{(s_{j},a_{j})}\min_{i=1,2}Q_{\varsigma_{i}}(s_{j},a_{j})}$ and $\omega\in\mathbb{R}^{+}$ is a normalization coefficient. We train BC-SAC for 400K gradient steps, collecting online data every 10 steps. We use the hyperparameters recommended in [lyu2024odrl].
H2O: H2O [niu2022trust] trains domain classifiers to estimate dynamics gaps and uses them as importance sampling weights during critic training. It also incorporates a CQL loss to encourage conservatism. Since the original H2O is designed for the Online-Offline setting, we adapt the objective to the Offline-Online setting. The critic loss is:
$\mathcal{L}_{\text{critic}}=$ $\ \mathbb{E}_{\mathcal{D}_{\text{on}}}\left[(Q_{\varsigma_{i}}(s,a)-y)^{2}\right]+\mathbb{E}_{\mathcal{D}_{\text{off}}}\left[\omega(s,a,s^{\prime})(Q_{\varsigma_{i}}(s,a)-y)^{2}\right]$
$+\beta_{\text{CQL}}\left(\mathbb{E}_{s\sim\mathcal{D}_{\text{off}},\ \hat{a}\sim\pi_{\varphi}(\cdot|s)}\left[\omega(s,a,s^{\prime})Q_{\varsigma_{i}}(s,\hat{a})\right]-\mathbb{E}_{\mathcal{D}_{\text{off}}}\left[\omega(s,a,s^{\prime})Q_{\varsigma_{i}}(s,a)\right]\right),\quad i\in\{1,2\},$ (42)
@@
BC-VGDF: BC-VGDF [xu2023cross] filters offline transitions whose estimated values align closely with those from the online environment. It trains an ensemble of dynamics models to predict next states from raw state-action pairs under the online dynamics. Each predicted next state is evaluated by the policy to obtain a value ensemble $\{Q(s^{\prime}_{i},a^{\prime}_{i})\}_{i=1}^{M}$, forming a Gaussian distribution. A fixed percentage ($\xi\%$) of offline transitions with the highest likelihood under this distribution are retained. The critic loss is:
$\mathcal{L}_{\text{critic}}=$ $\ \mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}_{\text{on}}}\left[\left(Q_{\varsigma_{i}}(s,a)-y\right)^{2}\right]$
$+\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}_{\text{off}}}\left[\mathbf{1}\left(\Lambda(s,a,s^{\prime})>\Lambda_{\xi\%}\right)\left(Q_{\varsigma_{i}}(s,a)-y\right)^{2}\right],\quad i\in\{1,2\},$ (43)